OpenStudy (anonymous):

The average hourly wage of employees of a certain company is $9.83. Assume the variable is normally distributed. If the standard deviation is $4.58, find the probability that a randomly selected employee earns less than $5.43

OpenStudy (anonymous):

The average hourly wage of employees of a certain company is $9.83. Assume the variable is normally distributed. If the standard deviation is $4.58, find the probability that a randomly selected employee earns less than $5.43. ------------------- Find the z-score of 5.43 z(5.43) = (5.43-9.83)/4.58 = -0.960699 ------- Then P(x < 5.43) = P(z < -0.960699) = 0.1684 = 16.84%